OpenStudy (anonymous):

Write x^2 - 6x + 7 = 0 in the form (x - a)^2 = b, where a and b are integers. (x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (anonymous):

@Whitemonsterbunny17 @ramit.dour @tester97

OpenStudy (anonymous):

(x - 3)^2 = 2

OpenStudy (anonymous):

\[(x-a)^2=b<=>x^2-2ax+a^2=b\] IDENTIFICATION:-2a=-6<=>a=3; \[x^2-6x+9=b<=>x^2-6x+9-b=0\] IDENTIFICATION:9-b=7<=>b=2; =>(x-3)^2=2

OpenStudy (anonymous):

Thank you! Can you help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A biologist created the following graph to show the relationship between the temperature of water (x), in degrees Celsius, and the number of insect larvae (y) in the water: graph of y equals minus 2 times the square of x plus 20 times x plus 400 What do the x-intercepts of the graph represent? There were 20 larvae at 0 degrees Celsius. There were 10 larvae at 0 degrees Celsius. The water has no larvae at -10 degrees Celsius and 20 degrees Celsius. The water has maximum larvae at -10 degrees Celsius and 20 degrees Celsius.

OpenStudy (anonymous):

Its C right?

OpenStudy (anonymous):

y=-2x^2+20x+400 x intercept results y=0; \[-2x^2+20x+400=0<=>\Delta=3600;x1=-10;x2=20\] 

OpenStudy (anonymous):

C is right

OpenStudy (anonymous):

Thank you!! I appreciate it!